# Advancing Reference-free Evaluation of Video Captions with Factual Analysis

### v1

@@ 6 Conclusion and Discussion
This work addresses the challenge of evaluating video captions across diverse domains without relying on human-annotated reference captions. We identified the limitations of existing metrics, and proposed a novel reference-free and factually grounded evaluation framework, VC-Inspector, based on LMMs. By delving into the factual analysis of captions, we empower the LMM with the factual grounding ability through instruction tuning, where we systematically create pseudo captions of diverse quality using an LLM. Experimental results across multiple domains demonstrate that VC-Inspector achieves high consistency with human evaluators, and outperforms existing metrics on detecting factual errors. Its versatility and interpretability make it a practical tool for evaluating the factual accuracy of video captions in real-world settings.
While our current focus is on factual errors, particularly object and action substitutions, the proposed paradigm can be potentially extended to simulate other types of errors (e.g., attribute replacements), as supported by our ablations on synthesizing only one type of error. Additionally, we recognize the importance of evaluating temporal coherence and narrative structure, which remain underexplored. Future work should expand this paradigm to simulate and detect a broader range of errors, enabling more comprehensive and context-aware evaluation strategies for video captions.
@@
The appendix is organized as follows. Section A lists the detailed prompts used in the proposed data generation pipeline, model fine-tuning, and explanation evaluation. Section B provides the evaluation of explanations generated by VC-Inspector, and section C presents additional visual examples with candidate captions of diverse quality.
Appendix A Prompts
The data generation prompts are reported on the following blocks:
  • Extract object – Prompt A
  • Extract action – Prompt A
  • Find similar object – Prompt A
  • Find similar action – Prompt A
  • Substitute object or action – Prompt A
The fine-tuning prompt and the prompt for evaluating the generated explanations are reported in Prompt A and Prompt A, respectively.
Prompt A.1: Prompt to extract object from a caption ### Instruction: Given the input text, generate a list of objects in the caption in the format of [ "Object1", "Object2", …]. Don’t include any verbs. ONLY REPLY THE ANSWER.
### Input: {{caption}}
### Output
Prompt A.2: Prompt to extract actions from a caption ### Instruction: Given the input text, generate a list of actions in the caption in the format of ["Action1", "Action2", …]. ONLY REPLY THE ANSWER.
### Input: {{caption}}
### Output
Prompt A.3: Prompt to find similar object given a object ### Instruction: Find the parent class of the given object and generate one of its child class that has a different meaning but shares the same parent. The new class cannot be a synonym or similar terms to the original object. It can be an antonym or any co-hyponym. For example, generate "dog" for "cat". ONLY REPLY THE NEW CLASS.
### Input: {{object}}
### Output:
Prompt A.4: Prompt to find similar action given an action ### Instruction: Find a different action that the subject can perform that has a different meaning than the input action. The new action cannot be a synonym or similar terms to the original action. For example, generate "put into" for "take out of". ONLY REPLY THE NEW ACTION.
### Input: {{action}}
### Output:
Prompt A.5: Prompt to substitute object or action given the caption and new object or actions ### Instruction: Substitute {{old_obj_act}} in {{caption}} as {{new_obj_act}} . Keep the answer in the same format as {{caption}} . ONLY REPLY THE ANSWER.
### Input: cap
### Output:
Prompt A.6: Fine-tuning prompt ### USER:
[[VIDEO]]
<caption> {{caption}} </caption> You are given a video and a caption describing the video content. Please rate the helpfulness, relevance, accuracy, level of details of the caption. The overall score should be on a scale of 1 to 5, where a higher score indicates better overall performance. Please first output a single line containing only one integer indicating the score. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias. STRICTLY FOLLOW THE FORMAT.
### ASSISTANT:
{{quality_score}}
The caption does not accurately capture the video content. For example, the actions ( {{wrong_act}} ) are incorrect / the objects ( {{wrong_obj}} ) are incorrect/the objects ( {{wrong_obj}} ) and actions ( {{wrong_act}} ) are incorrect.
Prompt A.7: Prompt for explanation evaluation [Context]
{{ground_truth_caption}}
[Caption]
{{caption_to_evaluate}}
[Groundtruth]
{{ground_truth_explanation}}
[End of Groundtruth]
[Assistant]
{{predicted_explanation}}
[End of Assistant]
[System]
We would like to request your feedback on the performance of an AI assistant in the response to the quality evaluation of the caption provided above with respect to a video. For your reference, the visual content in the video is represented with a few sentences describing the same video. You are also given a ground truth evaluation to that caption. Please rate the helpfulness, relevance, accuracy, level of details of the response by comparing to the ground truth and referring to the context information. Provide an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing the score. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias.
Appendix B Evaluation on Explanation Quality
@@
We further assess the quality of the explanations generated by VC-Inspector on two synthetic datasets, as the VATEX-Eval dataset does not offer ground truth explanations. Table 7 presents two metrics for this experiment. The BERT score measures the semantic similarity between the generated explanations and the pseudo ground truth explanations, according to pretrained textual embeddings. On the other hand, the LLM Score is derived by using the prompt detailed in section A to engage Llama-3-8B-Instruct in a LLM-as-a-judge approach, similar to [20]. We report the relative score assigned by the LLM that rescales the score received by the predicted explanations with respect to the score given to the ground truth explanations. Results show that the explanations provided by VC-Inspector, along with its quality assessment, align closely with the (pseudo) ground truth explanations obtained during the data generation process.
@@
Figure 5: Additional visual examples on VATEX-Eval. VC-Inspector identifies the incorrect objects (highlighted in red) and assigns scores aligned with human evaluators.
Appendix C Additional Visual Examples
Figure 5 presents additional visual examples of VC-Inspector applied to VATEX-Eval. Consistent with the results shown in Figure 4 of the main paper, VC-Inspector effectively identifies incorrect objects and actions in candidate captions, assigning quality scores that closely align with human judgments. In Figure 6, we showcase two video examples from ActivityNet-FG-Eval, each paired with candidate captions containing a progressively increasing number of factual inaccuracies. The model successfully detects these incorrect elements, provides detailed explanations, and yields scores that reflect the severity of the factual errors.

### v2

@@ 5 Conclusion
This work tackles the challenge of evaluating video captions across diverse domains without relying on human-annotated references. We conducted an extensive review of existing metrics, identified their limitations, and proposed VC-Inspector, a novel fact-aware evaluator for reference-free video caption evaluation. By incorporating factual analysis into caption assessment, we equipped a lightweight, open-source LMM with the factual grounding capabilities through training on captions of diverse quality, where pseudo captions with controllable factual errors were systematically generated. Experimental results across multiple domains demonstrate that VC-Inspector achieves high alignment with human judgments, and outperforms existing metrics in detecting factual errors. Its versatility and interpretability make it a practical tool for assessing factual accuracy in real-world video captioning scenarios. Furthermore, ablation studies on predicted explanations underscore its effectiveness in quality estimation and caption refinement.
Limitations
VC-Inspector primarily targets object and action correctness, which constitute major sources of factual errors in video captions. While these factors cover many common failure cases, other important aspects, e.g., attributes, spatial relationships, and fine-grained temporal ordering, are not explicitly modeled and remain directions for future work. Additionally, the training process partially relies on synthetically generated captions and pseudo-scores. Although these scores are deterministic and validated through correlation with human judgments, they may not fully capture the diversity of real-world captioning errors. Finally, while VC-Inspector is substantially more lightweight than proprietary LLM-based evaluators, it still depends on a multimodal backbone, resulting in higher computational costs compared to purely embedding-based metrics.
Despite these limitations, VC-Inspector represents a meaningful step toward explanation-aware, factually grounded, and reference-free video caption evaluation, providing a flexible foundation for future extensions.
@@
Appendix A Qualitative Results
Fig.˜5 presents visual examples of VC-Inspector outputs on the ActivityNet-FG-Eval (top) and VATEX-Eval (others) datasets. Without relying on reference captions, the model evaluates candidate captions based on the associated video content and produces quality scores that closely mirror human judgments. The explanation effectively pinpoints the factual inaccuracies, such as incorrect objects and/or actions in the candidate captions, and assigns scores accordingly.
In Fig.˜6, we showcase two video examples from ActivityNet-FG-Eval, each paired with candidate captions containing a progressively increasing number of factual errors. The model successfully detects these incorrect elements, provides detailed explanations, and yields scores that reflect the severity of the factual inaccuracies.
Appendix B Additional Results
B.1 Computational Efficiency
We assess computational efficiency by comparing the average runtime per video clip. EMScore (Shi et al., 2022), ViCLIPScore (Wang et al., 2024), and VC-Inspector-3B require 0.42, 0.34, and 0.30 seconds per clip, respectively, on a single A100 GPU. These results indicate that VC-Inspector achieves superior efficiency compared to existing methods. For EMScore and ViCLIPScore, we use the official implementations provided by the authors, while for VC-Inspector we employ vLLM (Kwon et al., 2023) with online serving. Computational efficiency is measured over 2,000 videos, each uniformly sampled into 16 frames.
Figure 5: Visual examples from ActivityNet-FG-Eval (top) and VATEX-Eval (others). VC-Inspector produces quality assessments consistent with ground truth scores, and explanatory insights into factual errors (highlighted in red).
B.2 Evaluation of Explanation Quality using Language Models
@@
We further assess the quality of the explanations generated by VC-Inspector on two synthetic datasets, as the VATEX-Eval dataset does not offer ground truth explanations. Table 7 presents two metrics for this experiment. The BERT score measures the semantic similarity between the generated explanations and the pseudo ground truth explanations, according to pretrained textual embeddings. On the other hand, the LLM Score is derived by using the prompt detailed in section D to engage Llama-3-8B-Instruct in a LLM-as-a-judge approach, similar to Liu et al. (2023a). We report the relative score assigned by the LLM that rescales the score received by the predicted explanations with respect to the score given to the ground truth explanations. Results show that the explanations provided by VC-Inspector, along with its quality assessment, align closely with the (pseudo) ground truth explanations obtained during the data generation process.
B.3 Complete Results for Caption Refinement
In §˜4.4, we present the experiment of caption refinement with CAPability benchmark, where we report the dimensions concerning objects and actions, as they are the primary focus of this work. We provide the F1 scores of all the 12 dimensions in Table˜8 for completeness.
@@ B.3 Complete Results for Caption Refinement
Feedback Provider
Task None Qwen2.5-VL-7B VC-Inspector-7B
Camera Angle 37.5 6.4 36.1
Camera Movement 26.4 0.6 29.3
Character Identification 29.6 17.3 30.5
Dynamic Object Num. 17.4 17.3 20.7
Event 58.2 18.8 58.6
Object Category 67.8 58.9 68.0
Object Color 73.0 51.9 72.3
Object Number 28.8 20.4 29.3
OCR 69.6 62.9 68.5
Scene 63.5 63.0 63.2
Spatial Relation 56.6 45.7 57.1
Style 84.3 78.7 84.2
Average 51.1 36.8 51.5
Table 8: F1 scores across different tasks. Video-based tasks are highlighted in blue, while image-based tasks are highlighted in beige. Best results are shown in bold.
Appendix C Training Hyperparameters
We report all training hyperparameters in Table˜9.
Hyperparameter Value
Training Configuration
Epochs 1
Global Batch Size 128
Learning Rate 1e-4
LR Scheduler Cosine (min: 1e-5)
Warmup Ratio 0.05
LoRA Configuration
Rank ($r$) 32
Alpha ($\alpha$) 32
Dropout 0.05
Module Freezing
Vision Encoder
Language Model
Projector
Video Processing
Number of Frames 32
Max. Pixels $224\times 224$
Table 9: Hyperparameters for instruction fine-tuning.
Appendix D Prompts
The data generation prompts are reported on the following blocks:
  • Extract object – Prompt D
  • Extract action – Prompt D
  • Find similar object – Prompt D
  • Find similar action – Prompt D
  • Substitute object or action – Prompt D
The data generation process provides the information of incorrect objects and actions. We format the explanations of these factual errors using the following template:
  • Captions without errors: The caption is helpful, relevant, accurate, and informative to the video content.
  • Captions with errors: The caption does not accurately capture the video content. For example, the actions ({{wrong_act}}) are incorrect / the objects ({{wrong_obj}}) are incorrect / the objects ({{wrong_obj}}) and actions ({{wrong_act}}) are incorrect.
These explanations are then used in Prompt D for training VC-Inspector.
The prompt for evaluating the generated explanations is in Prompt D. Prompt D is used to refine the caption based on the explanation.
Prompt D.1: Extract object from a caption ### Instruction: Given the input text, generate a list of objects in the caption in the format of [ "Object1", "Object2", ...]. Don’t include any verbs. ONLY REPLY THE ANSWER. ### Input: {{caption}} ### Output
Prompt D.2: Extract actions from a caption ### Instruction: Given the input text, generate a list of actions in the caption in the format of ["Action1", "Action2", ...]. ONLY REPLY THE ANSWER. ### Input: {{caption}} ### Output
Prompt D.3: Find similar object given a object ### Instruction: Find the parent class of the given object and generate one of its child class that has a different meaning but shares the same parent. The new class cannot be a synonym or similar terms to the original object. It can be an antonym or any co-hyponym. For example, generate "dog" for "cat". ONLY REPLY THE NEW CLASS. ### Input: {{object}} ### Output:
Prompt D.4: Prompt to find similar action given an action ### Instruction: Find a different action that the subject can perform that has a different meaning than the input action. The new action cannot be a synonym or similar terms to the original action. For example, generate "put into" for "take out of". ONLY REPLY THE NEW ACTION. ### Input: {{action}} ### Output:
Prompt D.5: Prompt to substitute object or action given the caption and new object or actions ### Instruction: Substitute {{old_obj_act} in {{caption}} as {{new_obj_act}}. Keep the answer in the same format as {{caption}}. ONLY REPLY THE ANSWER. ### Input: {{cap}} ### Output:
Prompt D.6: Fine-tuning prompt ### USER: {{VIDEO}} <caption>{{caption}}</caption> You are given a video and a caption describing the video content. Please rate the helpfulness, relevance, accuracy, level of details of the caption. The overall score should be on a scale of 1 to 5, where a higher score indicates better overall performance. Please first output a single line containing only one integer indicating the score. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias. STRICTLY FOLLOW THE FORMAT. ### ASSISTANT: {{quality_score}} {{explanation}}
Prompt D.7: Prompt for explanation evaluation [Context] {{ground_truth_caption}} [Caption] {{caption_to_evaluate}} [Groundtruth] {{ground_truth_explanation}} [End of Groundtruth] [Assistant] {{predicted_explanation}} [End of Assistant] [System] We would like to request your feedback on the performance of an AI assistant in the response to the quality evaluation of the caption provided above with respect to a video. For your reference, the visual content in the video is represented with a few sentences describing the same video. You are also given a ground truth evaluation to that caption. Please rate the helpfulness, relevance, accuracy, level of details of the response by comparing to the ground truth and referring to the context information. Provide an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. Please first output a single line containing the score. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias.
Prompt D.8: Prompt to refine caption based on explanation Previous Caption: {{caption}} Evaluation: {{explanation}} Based on the previous caption and evaluation, refine the caption to make it more accurate and detailed. Use the evaluation to find out the missing details and use the video frames to fix the missing details. Just output the refined caption, without any other text.
Appendix E Use of AI Assistance
During the preparation of this work, the authors utilized Cursor [^2] https://cursor.com for coding assistance and ChatGPT [^3] https://chatgpt.com/ for proofreading and language refinement. All input provided to these tools consisted of original content authored by the authors.